. Determine the angle of the two vectors a and b where their magnitudes are 11 units and 14 units, respectively. Their

resultant vector is 70.04.


1
Expert's answer
2021-10-25T09:57:07-0400

Given:

"a=11,\; b=14, \; |{\bf a}+{\bf b}|=7.04"


We have

"|{\bf a}+{\bf b}|^2=a^2+b^2-2ab\cos\theta"

"\cos\theta=\frac {a^2+b^2-|{\bf a}+{\bf b}|^2}{2ab}"

"\cos\theta=\frac {11^2+14^2-7.04^2}{2*11*14}=0.8683"

"\theta=29.7^{\circ}"